a truck of mass 1000 kg is moving with a constant vel. .......a force acts on it so that a retardation of 4 m/s2


is produced ..........................find the unbalanced force on the truck


PLS EXPLAIN!!!!!!!!!!!!!!!

Grade:9

1 Answers

Abhijat Pandey
12 years ago

Since retardation(-a) is happeing therefore, final vel. is zero and let initial vel. be u....

now, v=u+at

       (0=u-at) or (u=at)  ..................     (1)

we know that rate of change of momemtum is force applied.

therefore, (m2v2)- (m1v1)/t = F

 (where t= time in which truck comes in rest and F = unbalanced force applied,   also final momentum is 0)

now,

(0-mu)/t = F

from (1),

-mat/t = F

-ma = F

F = -ma 

F= -1000*4

F= -4000N

this is your answer!

plz approve!!